# Time Series Anomaly Detection in the Frequency Domain with Statistical Reliability

### v1

@@ D.1 Detailed descriptions of comparison methods
  • OC: In this method, we consider $p$-values conditioned on the process of the algorithm $\mathcal{A}$. The over-conditioned $p$-value is computed as
    $p_{k}^{\text{oc}}=\mathbb{P}_{\text{H}_{0},k}\left(Z\geq z^{\text{obs}}\,|\,Z\in\mathcal{Z}^{\text{oc}}\left(\bm{a}+\bm{b}z^{\text{obs}}\right)\right).$
    This method is computationally efficient, however, its power is low due to over-conditioning.
  • OptSeg-SI-oc (Duy et al., 2020): This method uses a $p$-value conditioned only on the process of dynamic programming algorithm. The over-conditioned $p$-value is computed as
    $p_{k}^{\text{OptSeg-SI-oc}}=\mathbb{P}_{\text{H}_{0},k}\left(Z\geq z^{\text{obs}}\,|\,Z\in\left\{r\in\mathbb{R}\,|\,\mathcal{S}_{\text{DP}}(r)=\mathcal{S}_{\text{DP}}\left(z^{\text{obs}}\right)\right\}\right).$
@@
    $p_{k}^{\text{naive}}=\mathbb{P}_{\text{H}_{0},k}\left(Z\geq z^{\text{obs}}\right).$
@@ D.2 Computational Time and Computer Resources
(a) Type I error rate
    ($M=512$)
(b) Type I error rate
    ($M=1024$)
(c) Power ($M=512$)
(d) Power ($M=1024$)
@@ D.3 Robustness of Type I Error Rate Control
Unknown noise variance. We generated 1000 null sequences in the same manner as the type I error rate experiment with known variance presented in Section 5.2. To estimate the variance $\sigma^{2}$ from the same data, we first applied CP candidate selection algorithm to identify the segments, and then computed the empirical variance of each segment for all frequencies. Since the estimated variance tended to be smaller than the true value, we adopted the maximum value for each frequency. Given $\hat{\sigma}_{f}$ as the average of the values, $\sigma$ could be estimated from the property of DFT as $\hat{\sigma}=\frac{\hat{\sigma}_{f}}{\sqrt{M}}$. The results for the significance levels $\alpha=0.01,0.05,0.1$ are shown in Figure 12 and the Proposed still could properly control the type I error rate.
Non-Gaussian noise. We considered the case where the noise followed the five non-Gaussian distributions:
@@
Correlated noise. We generated 10000 null sequences $\bm{x}=(x_{1},\dots,x_{N})^{\top}\sim\mathcal{N}(\bm{s},\Sigma)$, where mean vector $\bm{s}$ was specified in the same manner as described in the type I error rate experiment with known variance, and covariance matrix $\Sigma$ was defined as $\Sigma=\sigma^{2}\left(\rho^{|i-j|}\right)_{ij}\in\mathbb{R}^{N\times N}$ with $\sigma=1$ and $\rho\in\{0.025,0.05,0.075,0.1\}$, for $T=60$. After CP candidate selection, we conducted the hypothesis testing using the test statistic with $\sigma=1$. The results for the significance levels $\alpha=0.01,0.05,0.1$ are shown in Figure 14. For weak covariance, the Proposed could properly control the type I error rate. However, the type I error rate could not be controlled with increasing noise correlation. This remains a challenge for future work.
D.4 More Results on Real Data Experiments
Additional results for the signals of bearing 2, 3, and 4 before and after the anomaly of bearing 1 occurred are shown in Figures 15, 16, and 17. In panel (a) of each figure, the time variation of frequency spectra where CP candidate locations were falsely detected are shown for the period of 0.25–2.25 days when the frequency anomaly in bearing 1 did not actually exist. The results indicate that the inferences using $p$-values of the Proposed and OC are valid. In panel (b), the time variations of the BPFO harmonics where CP candidate locations were correctly detected are presented for the period of 4–6 days (bearing 2), 4.75–6.75 days (bearing 3), and 4–6 days (bearing 4), respectively. In these cases, although the detection was delayed by several days relative to the occurrence of the frequency anomaly in bearing 1, the outer race fault signatures were successfully identified in all bearings using the Proposed.
(a) $M=512$
(b) $M=1024$
@@ D.4 More Results on Real Data Experiments
(a) Inference on a falsely detected CP candidate location for 3260 Hz (around the 14th harmonic) on 0.25–2.25 days
(b) Inference on a truely detected CP candidate location for 1420 Hz (the 6th harmonic) on 4–6 days
@@
(a) Inference on a falsely detected CP candidate location for 3380 Hz (around the 14th harmonic) on 0.25–2.25 days
(b) Inferences on truely detected CP candidate locations for 1420 Hz (the 6th harmonic) on 4.75–6.75 days
@@
(a) Inference on a falsely detected CP candidate location for 3540 Hz (the 15th harmonic) on 0.25–2.25 days
(b) Inference on a truely detected CP candidate location for 1420 Hz (the 6th harmonic) on 4–6 days

### v2

@@ D.1 Detailed Descriptions of Comparison Methods
  • OC: In this method, we consider $p$-values conditioned on the process of the algorithm $\mathcal{A}$. The $p$-value is computed by obtaining the over-conditioned region $\mathcal{Z}^{\text{oc}}$ in (21) for the observed test statistic $T_{k}(\bm{x})$. This method is computationally efficient, however, its power is low due to over-conditioning.
  • OptSeg-SI-oc (Duy et al., 2020): This method uses a $p$-value conditioned only on the event $\mathcal{S}_{\text{DP}}$ in (21), that is, we consider conditioning on the process of dynamic programming algorithm.
@@
    $p_{k}^{\text{naive}}=\mathbb{P}_{\text{H}_{0},k}\left(T_{k}(\bm{X})\geq T_{k}(\bm{x})\right).$
@@ D.2 Computational Time and Computer Resources
(a) Type I error rate ($M=512$).
(b) Type I error rate ($M=1024$).
(c) Power ($M=512$).
(d) Power ($M=1024$).
@@ D.3 Robustness of Type I Error Rate Control
Unknown noise variance.
We generated 1000 null sequences in the same manner as the type I error rate experiment with known variance presented in Section 5.2. To estimate the variance $\sigma^{2}$ from the same data, we first applied CP candidate selection algorithm to identify the segments, and then computed the empirical variance of each segment for all frequencies. Since the estimated variance tended to be smaller than the true value, we adopted the maximum value for each frequency. Given $\hat{\sigma}_{f}$ as the average of the values, $\sigma$ could be estimated from the property of DFT as $\hat{\sigma}=\frac{\hat{\sigma}_{f}}{\sqrt{M}}$. The results for the significance levels $\alpha=0.01,0.05,0.1$ are shown in Figure 11 and the Proposed still could properly control the type I error rate.
Non-Gaussian noise.
We considered the case where the noise followed the five non-Gaussian distributions:
@@
Correlated noise.
We generated 10000 null sequences $\bm{x}=(x_{1},\dots,x_{N})^{\top}\sim\mathcal{N}(\bm{s},\Sigma)$, where mean vector $\bm{s}$ was specified in the same manner as described in the type I error rate experiment with known variance, and covariance matrix $\Sigma$ was defined as $\Sigma=\sigma^{2}\left(\rho^{|i-j|}\right)_{ij}\in\mathbb{R}^{N\times N}$ with $\sigma=1$ and $\rho\in\{0.025,0.05,0.075,0.1\}$, for $T=60$. After CP candidate selection, we conducted the hypothesis testing using the test statistic with $\sigma=1$. The results for the significance levels $\alpha=0.01,0.05,0.1$ are shown in Figure 13. For weak covariance, the Proposed could properly control the type I error rate. However, the type I error rate could not be controlled with increasing noise correlation. This remains a challenge for future work.
D.4 Sensitivity Study for the Penalty Parameter
This section presents a sensitivity study for the penalty parameter within the objective function in (5). As shown in Appendix A, the penalty parameter $\beta^{(d)}$ for each frequency $d$ can be theoretically derived using the BIC, leaving the penalty parameter $\gamma$ to be determined heuristically by users. Thus, we focus on the sensitivity study for $\gamma$ in this section. Specifically, we conducted the experiments for the type I error rate ($T=60$) and power ($\Delta=0.08$) on synthetic data as described in Section 5.2, while varying the hyper-parameter $\kappa$ in (20) across the values $\{1.0,1.5,2.0,2.5\}$. The results for the type I error rate and power experiments are shown in Figures 14 and 15, respectively. In Figure 14, the Proposed can control the type I error rate at the significance levels $\alpha=0.01,0.05,0.1$ for all values of $\kappa$. Additionally, Figure 15 shows that the power of the Proposed is the highest of all methods, regardless of the settings of $\kappa$. In particular, for the experiment with $M=512$, as the value of $\kappa$ increases, the power of the Proposed tends to be higher. This may be because a larger penalty makes distinct spectral changes more likely to be the subject of the hypothesis testing.
(a) $M=512$.
(b) $M=1024$.
@@ D.4 Sensitivity Study for the Penalty Parameter
(a) $M=512$.
(b) $M=1024$.
Figure 14: Sensitivity study of type I error rate control for the hyper-parameter $\kappa$ in the penalty parameter $\gamma$.
(a) $M=512$.
(b) $M=1024$.
Figure 15: Sensitivity study of power for the hyper-parameter $\kappa$ in the penalty parameter $\gamma$.
D.5 More Results on Real Data Experiments
Additional results for the signals of bearings 2, 3, and 4 before and after the spectral intensity enhancements occurred in the BPFO harmonics of bearing 1 are shown in Figures 16, 17, and 18. In panel (a) of each figure, the time variation of frequency spectra where CP candidate locations were falsely detected are shown for the period of 0.25–2.25 days when significant spectral changes in bearing 1 did not actually exist. The results indicate that the inferences using $p$-values of the Proposed and OC are valid. In panel (b), the time variations of the BPFO harmonics where CP candidate locations were correctly detected are presented for the period of 4–6 days (bearing 2), 4.75–6.75 days (bearing 3), and 4–6 days (bearing 4), respectively. In these cases, although the detection was delayed by several days relative to the occurrence of spectral amplification in bearing 1, the outer race fault signatures were successfully identified in all bearings using the Proposed.
(a) Inference on a falsely detected CP candidate location for 3260 Hz (around the 14th harmonic) on 0.25–2.25 days.
(b) Inference on a truely detected CP candidate location for 1420 Hz (the 6th harmonic) on 4–6 days.
@@ D.5 More Results on Real Data Experiments
(a) Inference on a falsely detected CP candidate location for 3380 Hz (around the 14th harmonic) on 0.25–2.25 days.
(b) Inferences on truely detected CP candidate locations for 1420 Hz (the 6th harmonic) on 4.75–6.75 days.
@@
(a) Inference on a falsely detected CP candidate location for 3540 Hz (the 15th harmonic) on 0.25–2.25 days.
(b) Inference on a truely detected CP candidate location for 1420 Hz (the 6th harmonic) on 4–6 days.